# System 0/1/2/3: Quad-process theory for multi-timescale embodied collective cognitive systems

### v2

@@ 1 Introduction
To understanding intelligence—whether biological or artificial—the diverse timescales at which cognition and adaptation occur must be considered. Existing models of cognition, such as dual-process theory, have significantly advanced our understanding of human intelligence by distinguishing between fast intuitive thinking (System 1) and slow deliberative reasoning (System 2) [69, 7]. However, these models remain limited in scope:
@@
Inside the brain, the free-energy principle proposed by Karl Friston provides a unifying framework for understanding adaptive behavior in biological systems [32, 101]. This principle suggests that intelligent agents strive to minimize surprise by continuously updating their internal world models [33, 137, 42, 44]. These world models are essential in both biological and artificial intelligence, enabling prediction, planning, and adaptive behavior.
However, cognition is not limited to the processes within an individual brain. Humans achieve long-term coordination and adaptation at the collective level by communicating, establishing rules, naming things, and creating agreements. These processes often rely on symbols, typically represented by language. In semiotics, it is widely recognized that symbol systems have arbitrariness, meaning that the relationship between the signifier and signified is not fixed and evolves across communities and time periods [17]. We emphasize this arbitrariness to highlight the core argument of this paper: the existence of super-slow dynamics in human (collective) intelligence. Human societies continually update shared symbol systems, i.e., systems of signs and their relationships with phenomena, over time through social interactions. This arbitrariness allows adaptive flexibility in semiotic communication. Humans utilize this semiotic plasticity, analogous to neural plasticity in the brain, to engage in intelligent activities that no single individual can achieve alone. We refer to this level as System 3 later in this paper.
Additionally, interactions between the brain, body, and environment demonstrate self-organizational and adaptive behavior, even without neural learning. For example, passive dynamic walkers can descend slopes without requiring computational control [88, 119]. These examples illustrate how body-environment interactions contribute to flexible, robust, and intelligent behavior generation. Although foundation models and multimodal LLMs are gaining attention [13, 109], robotics foundation models primarily focus on computational data processing and neglect embodied interactions [25, 2, 73, 23, 162, 70]. However, physical interactions that do not require neural processing exhibit super-fast dynamics. Recent AI studies have often overlooked this layer, which we refer to as System 0 in this paper.
The newly introduced terms of Systems 3 and 0 contrast with Systems 1 and 2 from dual-process theory [69]. By considering self-organizational dynamics both inside and outside the brain, we move beyond the computational view of intelligence to a self-organizational perspective.
@@ 2.1 Dual-process theory
Dual-process theory has been widely discussed across various domains, including cognitive science and behavioral economics [120, 27, 28, 69]. In the context of developing artificial cognitive systems that incorporate dual-process mechanisms, this section provides an overview of representative studies and discussions related to dual-process theory in AI and cognitive robotics.
@@ 2.1.1 Fast and Slow in AI and Cognitive systems
It has been argued that a combination of logical reasoning and deep learning is required to address these limitations. Bengio, [6] suggested the need to generalize from data to a consciousness stream consisting of a small number of sparsely connected concepts. Hybrid systems that integrate machine learning with symbolic reasoning, known as neuro-symbolic AI, have been actively investigated [117]. Symbolic AI is considered to correspond to System 2 and machine learning to System 1 [117]. Systems 1 and 2 are sometimes considered data-driven and knowledge-driven, respectively.
@@
Several frameworks have been proposed to implement System 2 using System 1. One of the most straightforward approaches is to use deep learning to acquire representations from the external world and then process these representations using symbolic AI. For symbolic AI to effectively manipulate these representations, it is necessary to acquire independent representations of each factor, a concept known in deep learning as disentangled representations [8]. Disentangled representations allow the system to clearly distinguish between different factors, making reasoning and manipulation more straightforward. For example, Higgins et al., [49] demonstrated that disentangled representations in generative models, such as the -VAE, enable a clear separation of independent factors of variation in data, thereby improving the ability to perform tasks such as causal reasoning and transfer learning. Locatello et al., [83] highlighted that although fully unsupervised disentanglement is challenging, leveraging some form of inductive bias or weak supervision can significantly improve the learning of disentangled representations.
@@
However, simply abstracting information from the external world may be insufficient. In practice, for an agent to adapt to a new environment, it must not only abstract the knowledge it has learned but also instantiate that knowledge within the new environment [14]. This requires the abilities to recognize similarities and differences between two environments and to determine what aspects should be forgotten during abstraction. Bengio’s “consciousness prior theory” suggests that such representation learning can be achieved by using conscious processing as a prior, which can be implemented through attention mechanisms [6]. Recurrent independent mechanisms (RIMs) introduce multiple recurrent neural network (RNN)-based modules that are sparsely connected based on attention mechanisms [86]. These modules operate independently and are activated only when relevant to the task, allowing for more efficient and specialized processing. In addition, a framework that generalizes abstracted knowledge as an autonomous skill that can be applied in new environments is necessary.Ellis et al., [26] proposed a method where programming languages are viewed as primitive skills and neural networks are used to guide the abstraction process. They demonstrated that this approach enables the automatic acquisition of skills applicable to new environments.
@@
To enable System 2-like reasoning, approaches such as chain-of-thought and tree-of-thought have been proposed [152, 158]. Rather than immediately providing an answer to a question all at once, these methods introduce a mechanism that explicitly outlines intermediate reasoning steps and verifies them incrementally until a conclusion is reached. Similar to how humans logically develop arguments in their minds, the model advances through multiple reasoning stages, making contradictions less likely to arise and thus making these approaches particularly suitable for tackling difficult mathematical problems as well as tasks involving long-text comprehension and summarization. Notably, simulating a tree-of-thought within prompts has been reported to enable LLMs to exhibit a more structured and deliberative inference process [84].
@@
Self-refine or self-reflection, first introduced as “Self-Refine” by [85], in LLMs can also be considered a method for enabling System 2-type reasoning. In this approach, the model itself examines whether its own output is correct or contains contradictions, and revises it as needed. Because the model first generates an answer and then reviews it, the process not only relies on intuition (System 1) but also incorporates deliberation and re-evaluation characteristic of System 2. A notable follow-up is Reflexion [118], which extends self-refine by letting LLMs revise their outputs multiple times. After each attempt, the model stores any mistakes or uncertainties and consults this record in future steps, thereby improving subsequent responses through repeated self-examination.
@@
In methods such as ReAct [159] and ToolFormer [114], LLMs acquire additional information or call external computational tools (e.g., Python, calculators, search engines) to improve their output. This is similar to human System 2 in that it implements a conscious, step-by-step process of not just generating answers intuitively (System 1), but also pausing to think about what information is lacking, gathering information from outside sources, and re-evaluating. ReAct is a prompting method that interleaves a LLM’s reasoning process with external tool usage, thus integrating the two. By combining introspective reasoning with external tool utilization, the model can actively think at each step while retrieving the necessary information, enabling it to solve more complex problems than would be possible with chain-of-thought alone. Toolformer is a model that, through self-supervised learning, acquires the ability to autonomously decide which external API tool to use and when to call upon it. This can be likened to how humans, when facing computations they cannot handle entirely in their head, may resort to writing on paper or consulting a dictionary. By allowing a pause in the continuous stream of text generation to explicitly call an API and wait for its results, the model incorporates a step-by-step, planful thinking process.
@@
Furthermore, the integration of reasoning with external tool usage has been actively advancing in the development of LLM agents in recent years [151]. For example, LangChain [18] serves as a framework for integrating LLMs with external tools, incorporating prompt designs and tool interactions inspired by approaches like ReAct. Additionally, efforts such as Auto-GPT [39] and BabyAGI [97] have emerged to develop agents that engage in continuous reasoning and planning while interacting with external environments to achieve long-term goals. Unlike single-turn Q&A models, these agents dynamically formulate plans, retrieve external resources as needed, and iteratively execute tasks to accomplish user-defined objectives. As a result, the "System 2"-like approach has become increasingly significant in the design and implementation of LLM agents.
@@ 2.1.2 Fast and Slow in Neuro-dynamical Systems
In the field of cognitive neurorobotics, Tani et al. developed various neuro-dynamical systems based on RNNs [133]. While their work primarily focused on the top-down intentional and bottom-up recognition processes underlying complex behavior and cognition in the brain, they also provided insights into how different levels of neural dynamics correspond to Systems 1 and 2.
@@
Their approaches were inspired by dynamical systems theory for autonomous agent behavior (both artificial and biological), emphasizing the tight coupling between an agent’s internal dynamics and the external dynamics of its environment [4, 5]. This original dynamical systems approach, which may correspond to Systems 0 and 1, explains an agent’s lower-level or reflexive behavior in response to environmental changes through the concept of entrainment into attractors. By combining this approach with the idea of forward models [155], which correspond with the so-called world models used in recent model-based reinforcement learning frameworks [43, 44, 45], early work on robotic navigation problems [131] demonstrated that the topological structure of the task space can be embedded as multiple attractors in the RNN by learning sensorimotor experiences.
@@
Tani’s subsequent research shifted towards understanding how higher-order cognitive functions are self-organized through the learning of continuous sensorimotor experiences, with minimal constraints imposed on neuro-dynamical systems and without introducing explicit mechanisms. Specifically, Tani et al. proposed hierarchical RNN models, such as the RNN with parametric bias (RNNPB) [132] and multiple timescale RNN (MTRNN) [156]. These models introduce a temporal hierarchy [71, 72], where the lower level exhibits a high temporal resolution, which may correspond to System 1, and the higher level exhibits a lower temporal resolution, which may correspond to System 2.
@@
In the RNNPB [132], the frequency of neural dynamics updates differs between levels. The lower level operates at the sensorimotor level with rapid updates, whereas the higher level updates more slowly. The higher level provides a vector called PB as a prior to the lower level, and the lower level generates sensorimotor predictions based on this PB from the higher level. The resulting prediction error, calculated as the difference between predicted and actual feedback, propagates upward to the higher level, updating its internal model of the world. This mutual interaction between the top-down predictive and bottom-up recognition processes can be understood as the RNN implementation of a predictive coding scheme [110]. Based on this RNNPB scheme, several attempts have been made to account for cognitive processes, such as object manipulation [66], imitation [66], language acquisition [126], and social interaction [20].
@@
Unlike RNNPB, which features different frequencies of neural dynamics updates, the MTRNN incorporates multiple timescales by assigning different time constants to each network level [156]. At a higher level, larger time constants produce slower neural dynamics, whereas at a lower level, smaller time constants lead to faster dynamics. This temporal constraint on each level allows the MTRNN to self-organize the so-called functional hierarchy, where the lower level with fast dynamics captures detailed primitives of sensorimotor experiences, and the higher level with slow dynamics captures abstract combinatorial representations of the primitives (see Figure 1). Several studies have demonstrated that primitives are acquired as limit cycles or fixed-point attractors at the lower level, whereas combinations of these primitives are represented as transient or chaotic dynamics at the higher level [156, 99, 98, 94]. In addition to cognitive neurorobotics studies, MTRNN-based models have been applied to more practical situations, such as flexible object manipulation [128] and door opening and entry with whole-body control [65], demonstrating their engineering utility.
@@
These fully deterministic neuro-dynamical systems, such as RNNPB and MTRNN, have been extended by incorporating uncertainty in both lower-level sensorimotor predictions (stochastic MTRNN; S-MTRNN) [95] and higher-level latent representations (predictive coding-inspired variational RNN; PV-RNN) [1], as well as in the multiple timescale recurrent state-space model (MTRSSM) [34]. The formulation of these models aligns with the free-energy principle [31, 32, 101]. These models demonstrate that noise at the lower level results in probabilistic transitions of primitives, whereas noise at the higher level produces similar phenomena.
@@
Models such as MTRNN, S-MTRNN, and PV-RNN, with their multiple timescales and complex temporal representations, have been explored not only in cognitive neurorobotics but also in the context of computational psychiatry [92]. Functional disconnections between hierarchies [157, 62] or aberrant estimation of uncertainty [60, 61] have been employed to explain pathologies in psychiatric disorders such as schizophrenia and autism spectrum disorder. These models offer a promising direction for understanding how neuro-dynamical processes contribute to cognitive dysfunction.

### v3

@@ 1 Introduction
To understanding intelligence—whether biological or artificial—the diverse timescales at which cognition and adaptation occur must be considered. Existing models of cognition, such as dual-process theory, have significantly advanced our understanding of human intelligence by distinguishing between fast intuitive thinking (System 1) and slow deliberative reasoning (System 2) (kahneman2011thinking; bengio2019system2). However, these models remain limited in scope:
@@
[…]
Inside the brain, the free-energy principle (FEP) introduced by Karl Friston offers a unifying account of how biological systems remain adaptive (friston2010free; Parr2022). According to FEP, an agent minimizes expected surprise by continually adjusting the parameters of a generative process that couples the organism to its niche. Whether one reads this process as an “internal model” or as a non-representational, action-oriented dynamic is left open—FEP itself is neutral on that point (Constant2021). What matters is that this temporally deep generative process supports prediction, planning, and other forms of adaptive behavior in both living organisms and artificial systems (friston2021world; taniguchi2023world).
However, cognition is not limited to the processes within an individual brain. Humans achieve long-term coordination and adaptation at the collective level by communicating, establishing rules, naming things, and creating agreements. These processes often rely on symbols, typically represented by language. In semiotics, it is widely recognized that symbol systems have arbitrariness, meaning that the relationship between the signifier and signified is not fixed and evolves across communities and time periods (Chandler2002). We emphasize this arbitrariness to highlight the core argument of this paper: the existence of super-slow dynamics in human (collective) intelligence. Human societies continually update shared symbol systems, i.e., systems of signs and their relationships with phenomena, over time through social interactions. This arbitrariness allows adaptive flexibility in semiotic communication. Humans utilize this semiotic plasticity, analogous to neural plasticity in the brain, to engage in intelligent activities that no single individual can achieve alone. We refer to this level as System 3 later in this paper.
Additionally, interactions between the brain, body, and environment demonstrate self-organizational and adaptive behavior, even without neural learning. For example, passive dynamic walkers can descend slopes without requiring computational control (mcgeer1990passive; shintake2018soft). These examples illustrate how body-environment interactions contribute to flexible, robust, and intelligent behavior generation. Although foundation models and multimodal LLMs are gaining attention (bommasani2021opportunities; radford2021learning), robotics foundation models primarily focus on computational data processing and neglect embodied interactions (driess2023palm; arai2024covla; kim2024openvla; dey2024revla; zhen20243d; Kawaharazuka16092024). However, physical interactions that do not require neural processing exhibit super-fast dynamics. Recent AI studies have often overlooked this layer, which we refer to as System 0 in this paper.
The newly introduced terms of Systems 3 and 0 contrast with Systems 1 and 2 from dual-process theory (kahneman2011thinking). By considering self-organizational dynamics both inside and outside the brain, we move beyond the computational view of intelligence to a self-organizational perspective.
@@
While the concept of System 3 is rooted in the Collective Predictive Coding (CPC) hypothesis introduced in Taniguchi (2024), this paper’s primary contribution is to extend the traditional System 1/2 framework into a comprehensive quad-process model. The key novelties are: first, the introduction of System 0 to account for pre-cognitive embodied processes, thereby completing the 0/1/2/3 architecture; and second, the grounding of this entire multi-timescale framework in an extended interpretation of Bergson’s philosophy of time, as detailed in 4.
@@ 2.1 Dual-process theory
Dual-process theory has been widely discussed across various domains, including cognitive science and behavioral economics (sloman1996empirical; evans2008dual; evans2013dual; kahneman2011thinking). In the context of developing artificial cognitive systems that incorporate dual-process mechanisms, this section provides an overview of representative studies and discussions related to dual-process theory in AI and cognitive robotics.
@@ 2.1.1 Fast and Slow in AI and Cognitive systems
It has been argued that a combination of logical reasoning and deep learning is required to address these limitations. bengio2017consciousness suggested the need to generalize from data to a consciousness stream consisting of a small number of sparsely connected concepts. Hybrid systems that integrate machine learning with symbolic reasoning, known as neuro-symbolic AI, have been actively investigated (sheth2023neurosymbolic). Symbolic AI is considered to correspond to System 2 and machine learning to System 1 (sheth2023neurosymbolic). Systems 1 and 2 are sometimes considered data-driven and knowledge-driven, respectively.
@@
Several frameworks have been proposed to implement System 2 using System 1. One of the most straightforward approaches is to use deep learning to acquire representations from the external world and then process these representations using symbolic AI. For symbolic AI to effectively manipulate these representations, it is necessary to acquire independent representations of each factor, a concept known in deep learning as disentangled representations (bengio2013representation). Disentangled representations allow the system to clearly distinguish between different factors, making reasoning and manipulation more straightforward. For example, higgins2017beta demonstrated that disentangled representations in generative models, such as the -VAE, enable a clear separation of independent factors of variation in data, thereby improving the ability to perform tasks such as causal reasoning and transfer learning. locatello2019challenging highlighted that although fully unsupervised disentanglement is challenging, leveraging some form of inductive bias or weak supervision can significantly improve the learning of disentangled representations.
@@
However, simply abstracting information from the external world may be insufficient. In practice, for an agent to adapt to a new environment, it must not only abstract the knowledge it has learned but also instantiate that knowledge within the new environment (booch2021thinking). This requires the abilities to recognize similarities and differences between two environments and to determine what aspects should be forgotten during abstraction. Bengio’s “consciousness prior theory” suggests that such representation learning can be achieved by using conscious processing as a prior, which can be implemented through attention mechanisms (bengio2017consciousness). Recurrent independent mechanisms (RIMs) introduce multiple recurrent neural network (RNN)-based modules that are sparsely connected based on attention mechanisms (madan2021fast). These modules operate independently and are activated only when relevant to the task, allowing for more efficient and specialized processing. In addition, a framework that generalizes abstracted knowledge as an autonomous skill that can be applied in new environments is necessary.ellis2021dreamcoder proposed a method where programming languages are viewed as primitive skills and neural networks are used to guide the abstraction process. They demonstrated that this approach enables the automatic acquisition of skills applicable to new environments.
@@
To enable System 2-like reasoning, approaches such as chain-of-thought and tree-of-thought have been proposed (wei2022chain; yao2024tree). Rather than immediately providing an answer to a question all at once, these methods introduce a mechanism that explicitly outlines intermediate reasoning steps and verifies them incrementally until a conclusion is reached. Similar to how humans logically develop arguments in their minds, the model advances through multiple reasoning stages, making contradictions less likely to arise and thus making these approaches particularly suitable for tackling difficult mathematical problems as well as tasks involving long-text comprehension and summarization. Notably, simulating a tree-of-thought within prompts has been reported to enable LLMs to exhibit a more structured and deliberative inference process (long2023large).
@@
Self-refine or self-reflection, first introduced as “Self-Refine” by (madaan2023self), in LLMs can also be considered a method for enabling System 2-type reasoning. In this approach, the model itself examines whether its own output is correct or contains contradictions, and revises it as needed. Because the model first generates an answer and then reviews it, the process not only relies on intuition (System 1) but also incorporates deliberation and re-evaluation characteristic of System 2. A notable follow-up is Reflexion (shinn2023reflexion), which extends self-refine by letting LLMs revise their outputs multiple times. After each attempt, the model stores any mistakes or uncertainties and consults this record in future steps, thereby improving subsequent responses through repeated self-examination.
@@
In methods such as ReAct (yao2023react) and ToolFormer (schick2023toolformer), LLMs acquire additional information or call external computational tools (e.g., Python, calculators, search engines) to improve their output. This is similar to human System 2 in that it implements a conscious, step-by-step process of not just generating answers intuitively (System 1), but also pausing to think about what information is lacking, gathering information from outside sources, and re-evaluating. ReAct is a prompting method that interleaves a LLM’s reasoning process with external tool usage, thus integrating the two. By combining introspective reasoning with external tool utilization, the model can actively think at each step while retrieving the necessary information, enabling it to solve more complex problems than would be possible with chain-of-thought alone. Toolformer is a model that, through self-supervised learning, acquires the ability to autonomously decide which external API tool to use and when to call upon it. This can be likened to how humans, when facing computations they cannot handle entirely in their head, may resort to writing on paper or consulting a dictionary. By allowing a pause in the continuous stream of text generation to explicitly call an API and wait for its results, the model incorporates a step-by-step, planful thinking process.
@@
Furthermore, the integration of reasoning with external tool usage has been actively advancing in the development of LLM agents in recent years (wang2024survey). For example, LangChain (langchain) serves as a framework for integrating LLMs with external tools, incorporating prompt designs and tool interactions inspired by approaches like ReAct. Additionally, efforts such as Auto-GPT (autogpt2023) and BabyAGI (babyagi2023) have emerged to develop agents that engage in continuous reasoning and planning while interacting with external environments to achieve long-term goals. Unlike single-turn Q&A models, these agents dynamically formulate plans, retrieve external resources as needed, and iteratively execute tasks to accomplish user-defined objectives. As a result, the "System 2"-like approach has become increasingly significant in the design and implementation of LLM agents.
@@ 2.1.2 Fast and Slow in Neuro-dynamical Systems
In the field of cognitive neurorobotics, Tani et al. developed various neuro-dynamical systems based on RNNs (Tani2016). While their work primarily focused on the top-down intentional and bottom-up recognition processes underlying complex behavior and cognition in the brain, they also provided insights into how different levels of neural dynamics correspond to Systems 1 and 2.
@@
Their approaches were inspired by dynamical systems theory for autonomous agent behavior (both artificial and biological), emphasizing the tight coupling between an agent’s internal dynamics and the external dynamics of its environment (Beer1995; Beer1996). This original dynamical systems approach, which may correspond to Systems 0 and 1, explains an agent’s lower-level or reflexive behavior in response to environmental changes through the concept of entrainment into attractors. By combining this approach with the idea of forward models (Wolpert1995), which correspond with the so-called world models used in recent model-based reinforcement learning frameworks (Ha2018b; Hafner2018; Hafner2020a), early work on robotic navigation problems (Tani1996) demonstrated that the topological structure of the task space can be embedded as multiple attractors in the RNN by learning sensorimotor experiences.
@@
Tani’s subsequent research shifted towards understanding how higher-order cognitive functions are self-organized through the learning of continuous sensorimotor experiences, with minimal constraints imposed on neuro-dynamical systems and without introducing explicit mechanisms. Specifically, Tani et al. proposed hierarchical RNN models, such as the RNN with parametric bias (RNNPB) (Tani2003) and multiple timescale RNN (MTRNN) (Yamashita2008). These models introduce a temporal hierarchy (Kiebel2008; Kiebel2009), where the lower level exhibits a high temporal resolution, which may correspond to System 1, and the higher level exhibits a lower temporal resolution, which may correspond to System 2.
@@
In the RNNPB (Tani2003), the frequency of neural dynamics updates differs between levels. The lower level operates at the sensorimotor level with rapid updates, whereas the higher level updates more slowly. The higher level provides a vector called PB as a prior to the lower level, and the lower level generates sensorimotor predictions based on this PB from the higher level. The resulting prediction error, calculated as the difference between predicted and actual feedback, propagates upward to the higher level, updating its internal model of the world. This mutual interaction between the top-down predictive and bottom-up recognition processes can be understood as the RNN implementation of a predictive coding scheme (Rao1999). Based on this RNNPB scheme, several attempts have been made to account for cognitive processes, such as object manipulation (Ito2006a), imitation (Ito2006a), language acquisition (Sugita2005), and social interaction (Chen2016).
@@
Unlike RNNPB, which features different frequencies of neural dynamics updates, the MTRNN incorporates multiple timescales by assigning different time constants to each network level (Yamashita2008). At a higher level, larger time constants produce slower neural dynamics, whereas at a lower level, smaller time constants lead to faster dynamics. This temporal constraint on each level allows the MTRNN to self-organize the so-called functional hierarchy, where the lower level with fast dynamics captures detailed primitives of sensorimotor experiences, and the higher level with slow dynamics captures abstract combinatorial representations of the primitives (see Figure 1). Several studies have demonstrated that primitives are acquired as limit cycles or fixed-point attractors at the lower level, whereas combinations of these primitives are represented as transient or chaotic dynamics at the higher level (Yamashita2008; Nishimoto2009; Namikawa2011; Murata2018). In addition to cognitive neurorobotics studies, MTRNN-based models have been applied to more practical situations, such as flexible object manipulation (Suzuki2018) and door opening and entry with whole-body control (Ito2022), demonstrating their engineering utility.
@@
These fully deterministic neuro-dynamical systems, such as RNNPB and MTRNN, have been extended by incorporating uncertainty in both lower-level sensorimotor predictions (stochastic MTRNN; S-MTRNN) (Murata2015a) and higher-level latent representations (predictive coding-inspired variational RNN; PV-RNN) (Ahmadi2019), as well as in the multiple timescale recurrent state-space model (MTRSSM) (Fujii2023). The formulation of these models aligns with the free-energy principle (Friston2005; friston2010free; Parr2022). These models demonstrate that noise at the lower level results in probabilistic transitions of primitives, whereas noise at the higher level produces similar phenomena.
@@
Models such as MTRNN, S-MTRNN, and PV-RNN, with their multiple timescales and complex temporal representations, have been explored not only in cognitive neurorobotics but also in the context of computational psychiatry (Montague2012). Functional disconnections between hierarchies (Yamashita2012a; Idei2021a) or aberrant estimation of uncertainty (Idei2018; Idei2020) have been employed to explain pathologies in psychiatric disorders such as schizophrenia and autism spectrum disorder. These models offer a promising direction for understanding how neuro-dynamical processes contribute to cognitive dysfunction.